A tennis court in the shape of a rectangle has an area of 7200 square feet. One pair of sides measures twice the length of the other pair of sides. What are the dimensions of the tennis court?
Answer by rfer(16322)

2x*x=7200
2x^2=7200
x^2=7200/2
x=sqrt3600
x=60 ft
2x=120 ft